OpenStudy (anonymous):

What is the x-intercept in this linear equation? 2y - 4x = -12

OpenStudy (anonymous):

Take the 4x and do the following -4x = -12 divide -12 by -4x to get x = 3

OpenStudy (anonymous):

2y - 4x = -12 write slope intercept y=(4x-12)\2 y= 2x-6 0=2x-6 6=2x x=6\2 x=3

OpenStudy (anonymous):

It is positive this time because both the 4x and 12 were negative

OpenStudy (anonymous):

wat i dont get it sorry

OpenStudy (anonymous):

take the 4x and 12 and divide the 12 by the 4 to get 3 therefore x = 3

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

Your welcome! :D